3. If the barometric pressure is 29.75 in Hg,convert (a) 27 psia to kPag, b) 24 in Hg vacuum to psia, c) 12

psia to mm Hg vacuum, d) 14 in Hg gage to kPaa.


1
Expert's answer
2021-09-07T23:52:23-0400

a)

1psia=6.9 kPaa

27psia=27*6.9kPaa

=186.3kPaa

kPag=kPaa-101.15

Hence;

12psia=(186.3kpaa-101.15)kPag

=85.15kPag

b)

2.03 in Hg vacuum =1 psia

Therefore;

24 in Hg vacuum="\frac{24}{2.03}"

=11.823psia

c)

1psia=2.03 in Hg vacuum

Therefore, 12psia= 12*2.03 in Hg vacuum

=24.36in Hg vacuum

1 in Hg vacuum =25.4mm Hg vacuum

Hence,

=24.36*25.4

=618.744mm Hg vacuum

d)

1.02 in Hg gauge= 97.7 kPaa

14 in Hg gauge= "\frac{14*97.7}{1.02}"

=1340.98kPaa